# On Dirichlet Series Involving $\zeta(s)$ and Extensions of the Euler–Mascheroni Constant

### v1

@@ Abstract.
In this paper, we introduce a class of Dirichlet series defined in terms of the Riemann zeta-function, motivated by the study of their special values, and establish integral representations for these series. We also define an extension of the Euler–Mascheroni constant and express certain special values explicitly in terms of the Bendersky constants. These results provide a unified framework for evaluating Dirichlet series involving the Riemann zeta-function at integer arguments, together with the associated number-theoretic constants.
@@ 1. Introduction
As usual, $\zeta(s)$ denotes the Riemann zeta function. Throughout this paper, we denote by $\mathbb{N}=\{1,2,3,\dots\}$ the set of natural numbers and by $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$ the set of nonnegative integers. The main object of this paper is the generating function of $\zeta(s)$ defined by the following Dirichlet series,
(1.1) ${D}(s_{1},s_{2};\lambda):=\sum\limits_{k=1}^{\infty}\dfrac{\zeta(s_{1}+k)-1}{(k+\lambda)^{s_{2}}}.$
Here $s_{1},s_{2}\in\mathbb{C}$ with $s_{1}\notin\mathbb{Z}_{\leq 0}$, and $\lambda\in\mathbb{C}$ with $\lambda\notin\mathbb{Z}_{<0}$. To investigate the special values of ${D}{}(s_{1},s_{2};\lambda)$, we introduce the following function, which serves as a natural generalization of the Euler constant.
(1.2) $\gamma(\alpha,\beta;\mu):=\lim\limits_{n\to\infty}\left\{\sum\limits_{l=1}^{\mu}\dfrac{1}{l^{\beta}}\sum\limits_{k=1}^{n}\dfrac{1}{k^{\alpha+l-\mu}}-\sum\limits_{k=2}^{n}{\rm{Li}}_{\beta}\left(\frac{1}{k}\right)k^{\mu-\alpha}\right\}.$
Here $\alpha,\beta\in\mathbb{C}$ with ${\Re(\alpha)}>0$ and $\mu\in\mathbb{N}_{0}$, and ${\rm{Li}}_{s}(z)=\sum\limits_{k=1}^{\infty}z^{k}/k^{s}$ denotes the polylogarithm; when $\mu=0$, the corresponding sum is omitted. For $\mu\geq 1$, the following equality holds:
(1.3) $D(\alpha,\beta;\mu)=\sum\limits_{l=1}^{\mu}\dfrac{1}{l^{\beta}}-\gamma(\alpha,\beta;\mu).$
This paper establishes fundamental properties of the function $D(s_{1},s_{2};\lambda)$ introduced above, including integral representations (Theorem 2.1 and Lemma 4.1). It further provides closed-form expressions for its special values (Theorems 2.2–2.4 and 2.6) in terms of values of the Riemann zeta function and certain classical arithmetic constants. These results form a unified framework for $D(s_{1},s_{2};\lambda)$ that integrates and acknowledges prior studies and clarifies structural features that have not been previously made explicit.
Historically, the origin of the Dirichlet series $(1.1)$ can be traced back to Goldbach’s classical theorem, which was communicated in a letter to D.  Bernoulli. In terms of the Riemann zeta function $\zeta(s)$, Goldbach’s theorem takes the following form (see for example [Sr2, Chapter 1]),
(1.4) $\sum_{k=2}^{\infty}\left\{\zeta(k)-1\right\}=1,$
which corresponds to the case D(1,0;1)=1{D}(1,0;1)=1 in $(1.3)$. For the case $s_{1}=1,s_{2}=1,\lambda=1$, the identity $(1.3)$ coincides with Euler’s classical formula (cf. [Eu]),
@@
Following these pioneering contributions, several further classical results were obtained in this direction (see, e.g., [Gl], [Gl2], [Raj], [Ras], [Ap]). A comprehensive historical overview of formulas expressing infinite series involving the Riemann zeta function in closed form can be found in [Sr] and [SC, Chapter 3]. Building upon these historical developments, it is then natural to explore explicit formulas for the special values of $D(s_{1},s_{2};\lambda)$. In the present paper, we base our investigation on the formulas due to Srivastava [Sr, (5.3)] (see also [SC, 3.4.(500)]), Adamchik–Srivastava [AS, Proposition 4] (see also [SC, Proposition 3.8]), and Choi–Srivastava [CS, (4.40)] (see also [SC, 3.4.(543)]), which will be recalled in Section 2.
@@ 2. Main Results
We use the following notation throughout this paper. As usual, $\Gamma(s)$ and $\zeta(s,\alpha)$ denote the gamma function and the Hurwitz zeta function, respectively. The symbol $\int_{+\infty}^{(0+)}$ denotes integration along a Hankel contour, which starts at positive infinity on the real axis, encircles the origin with a small radius in the positive (counterclockwise) direction, and then returns to the starting point. We denote the generalized harmonic numbers by $H_{m}^{(n)}:=\sum_{k=1}^{m}k^{-n}$, with $H_{m}:=H_{m}^{(1)}$ denoting the ordinary harmonic numbers, and we set $H_{0}^{(n)}:=0$. We also write $B_{j}$ for the $j$th Bernoulli number. As usual, the binomial coefficient $\binom{x}{k}$ is defined by $x(x-1)\cdots(x-k+1)/k!$ for $k\in\mathbb{N}_{0}$.
@@
Let $\lambda\in\mathbb{C}$ with $\lambda\notin\mathbb{Z}_{<0}$. The series ${D}(s_{1},s_{2};\lambda)$ in $(1.1)$ converges absolutely for any $s_{1},s_{2}\in\mathbb{C}$ except when $s_{1}\in\{0,-1,-2,\dots\}$. Furthermore, the following integral representation holds:
(2.1) ${D}(s_{1},s_{2};\lambda)=\dfrac{e^{-\pi is_{2}}\Gamma(1-s_{2})}{2\pi i(e^{2\pi is_{1}}-1)}\int\limits_{+\infty}^{(0+)}\int\limits_{+\infty}^{(0+)}\dfrac{v^{s_{1}-1}w^{s_{2}-1}}{e^{v+\lambda w}(e^{v}-1)}\sum\limits_{k=1}^{\infty}\dfrac{(v/e^{w})^{k}}{\Gamma(s_{1}+k)}dvdw.$
This representation also defines a holomorphic function on the entire $(s_{1},s_{2})$-plane except when $s_{1}\in\{0,-1,-2,\dots\}$.
@@
Let $\gamma(\alpha,\beta;\mu)$ be the generalized Euler constant defined in $(1.2)$. For $n\in\mathbb{N}$ and $m,\mu\in\mathbb{N}_{0}$,
@@
Since classical times, it has been recognized that the arithmetic nature of Euler’s constant remains unresolved: its transcendence and any closed-form representation are unknown, and even its irrationality has not been established (for a comprehensive survey, see [Rag]). Likewise, a closed-form expression for $\gamma(n,m;\mu)$ $(n,m\in\mathbb{N},\mu\in\mathbb{N}_{0})$ also appears to be intractable. As shown in equation $(1.5)$, the classical case $D(1,1;1)$ was already evaluated by Euler [Eu]. Furthermore, for $\mu=2,3$, the corresponding expressions $D{}(1,1;\mu)$ were given by Srivastava [Sr, (5.3)] and Choi–Srivastava [CS, (4.40)], in terms of the Stirling and Glaisher–Kinkelin constants, respectively:
[…]
Here $A_{0}=\sqrt{2\pi}$ is the Stirling constant; $A_{1}$ is the Glaisher–Kinkelin constant, defined as follows:
$\log A_{0}=\log\sqrt{2\pi}=\lim\limits_{n\rightarrow\infty}\left\{\sum\limits_{k=1}^{n}\log k-\left({n}+\frac{1}{2}\right)\log n+n\right\},$
$\log A_{1}=\lim\limits_{n\rightarrow\infty}\left\{\sum\limits_{k=1}^{n}k\log k-\left(\ \frac{n^{2}}{2}+\frac{n}{2}+\frac{1}{12}\right)\log n+\frac{n^{2}}{4}\right\}.$
The Glaisher–Kinkelin constant and its generalizations originally arose in the study of the hyperfactorial function and its extensions. Generally, the constants $A_{m}$ ($m\in\mathbb{N}_{\geq 2}$) are referred to as the Bendersky constants (cf. [Be]; see also [Ad, Section 5], [Co], and [PG]). They are defined via the following limit, and will be discussed further in Section 5.
$\log A_{m}:=\lim\limits_{n\to\infty}\left\{\sum\limits_{k=1}^{n}k^{m}\log k-p(n,m)\right\},$
@@
(2.3) $\begin{array}[]{ll}&p(n,m)=\left(\dfrac{n^{m+1}}{m+1}+\dfrac{n^{m}}{2}\right)\log n-\dfrac{n^{m+1}}{(m+1)^{2}}\\[10.0pt] &\quad+\;m!\displaystyle\sum\limits_{r=1}^{m}\dfrac{B_{r+1}}{(r+1)!}\dfrac{n^{m-r}}{(m-r)!}\left(\log n+(1-\delta_{m,r})\displaystyle\sum\limits_{i=1}^{r}\dfrac{1}{m-i+1}\right),\end{array}$
where $\delta_{m,r}$ denotes the Kronecker delta. When $m=0$, the summation over $r$ is understood to be omitted. With these preliminaries in place, we are now ready to state one of our main theorems.
Theorem 2.3 (A relation between the generalized Euler constant and the Bendersky constants).
@@
(2.4) $\gamma(1,1;\mu)=\frac{1}{\mu}\gamma+\sum\limits_{j=0}^{\mu-2}\binom{\mu-1}{j}\log A_{j},$
namely,
$D(1,1;\mu)=H_{\mu}-\frac{1}{\mu}\gamma-\sum\limits_{j=0}^{\mu-2}\binom{\mu-1}{j}\log A_{j}.$
Here, the summation over $j$ is understood to be empty when $\mu=1$.
@@
$\log A_{-1}(a):=\lim\limits_{n\to\infty}\left\{\sum\limits_{k=1}^{n}\frac{1}{k+a}\log k-\frac{1}{2}\log^{2}n\right\}.$
It follows that
$D(1,1;0)=\sum\limits_{l=1}^{\infty}(-1)^{l}\zeta^{\prime}(l+1)=\log A_{-1}(0)-\log A_{-1}(1).$
Taking the exponential of both sides, we obtain equivalently
$A_{-1}(a)=\lim\limits_{n\to\infty}\prod\limits_{k=1}^{n}k^{\frac{1}{k+a}}/n^{\log\sqrt{n}},\quad e^{D(1,1;0)}=\prod\limits_{n=1}^{\infty}n^{\frac{1}{n}-\frac{1}{n+1}}.$
In the above theorem, $\log A_{-1}(0)$ coincides with the first Stieltjes constant $\gamma_{1}$ (see, e.g., [Er, 1.12 (17)]). Theorems 2.3 and 2.4 indicate that the case $m=1$ is exceptional, while obtaining comparably simple closed forms for $D(n,m;\mu)$ with $m\geq 2$ is in general a difficult problem, reminiscent of the situation for $\zeta(s)$ at odd integer points $s\geq 3$. Nevertheless, it is possible to derive the following representations for the corresponding series.
@@
$\sum\limits_{m=0}^{\infty}D(n,m;\mu)=n+H_{\mu-1}-\gamma(n,1;\mu-1)-\sum\limits_{k=2}^{n}\zeta(k),$
$\sum\limits_{m=0}^{\infty}(-1)^{m}D(n,m;\mu)=n-H_{\mu+1}+\gamma(n,1;\mu+1)-\sum\limits_{k=2}^{n}\zeta(k),$
$\sum\limits_{m=0}^{\infty}D(n,2m+1;\mu)=H_{\mu}-\dfrac{1}{2}\Big\{\gamma(n,1;\mu+1)+\gamma(n,1;\mu-1)+\dfrac{1}{\mu(\mu+1)}\Big\}.$
@@
We now turn to the case $D(n,-m;\mu)$. In the fundamental case $n=\mu=1$, Adamchik and Srivastava first obtained the following explicit formula [AS, Proposition 4]; see also [SC, Proposition 3.8].
@@
Here $S(m,k)$ denotes the Stirling number of the second kind, defined by
@@
Building on this established result, we obtain the following theorem for $D(n,-m;\mu)$, which reduces to the case $n=\mu=1$.
@@
$D(n,-m;\mu)=\displaystyle\sum\limits_{l=0}^{m}\dbinom{m}{l}(\mu-n)^{m-l}D(1,-l;1)-\displaystyle\sum\limits_{k=2}^{n}\left\{\zeta(k)-1\right\}(k+\mu-n)^{m}.$
We also note that the following index-shifting relation for $D(n,\pm m;\mu)$ follows directly from Definition $(1.1)$. That is, for $n,\mu\in\mathbb{N}$ and $m\in\mathbb{Z}\setminus\{0\}$,
[…]
By combining the above index-shifting relations with Theorems 2.3 and 2.4, we obtain several structural relations among the values of $D(n,1;\mu)$.
As a concluding result of this section, we establish a functional relation for $D(s_{1},s_{2};\lambda)$ in the variable $s_{2}$. In particular, for D(n,m;0)D(n,-m;0) with $n,m\in\mathbb{N}$ and $n>1$, this relation provides a representation of these values.
@@
Let $s_{1},s_{2}\in\mathbb{C}$ with $s_{1}\notin\mathbb{Z}_{\leq 0}$. Assume that $\Re(s_{1}-\lambda)>1$, $\Re(s_{2})<0$, and $1<\lambda\leq 0$. Then the following functional relation holds:
$D(s_{1},s_{2};\lambda)=\sum_{\begin{subarray}{c}k\geq 2,\;l\in\mathbb{Z}\end{subarray}}\frac{e^{2\pi il\lambda}\,\Gamma(1-s_{2})}{k^{\,s_{1}-\lambda}\,(2\pi il-\log k)^{\,1-s_{2}}}.$

### v2

@@ Abstract.
In this paper, we introduce a class of Dirichlet series defined in terms of the Riemann zeta-function and establish integral representations for these series. Motivated by the study of their special values, we define an extension of the Euler–Mascheroni constant and obtain explicit expressions for certain values in terms of the Bendersky constants. These results provide a unified framework for the evaluation of Dirichlet series involving the Riemann zeta-function at integer arguments, together with the associated number-theoretic constants.
@@ 1. Introduction
As usual, $\zeta(s)$ denotes the Riemann zeta function. Throughout this paper, we denote by $\mathbb{N}=\{1,2,3,\dots\}$ the set of natural numbers and by $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$ the set of nonnegative integers. The main object of this paper is the Dirichlet series
(1.1) $D(s_{1},s_{2};\lambda):=\sum_{k=1}^{\infty}\frac{\zeta(s_{1}+k)-1}{(k+\lambda)^{s_{2}}},$
which may be regarded as a generating function associated with the Riemann zeta function. Here $s_{1},s_{2}\in\mathbb{C}$ with $s_{1}\notin\mathbb{Z}_{\leq 0}$, and $\lambda\in\mathbb{C}$ with $\lambda\notin\mathbb{Z}_{<0}$. To investigate the special values of $D(s_{1},s_{2};\lambda)$, we introduce the function
(1.2) $\gamma(\alpha,\beta;\mu):=\lim_{n\to\infty}\left\{\sum_{l=1}^{\mu}\frac{1}{l^{\beta}}\sum_{k=1}^{n}\frac{1}{k^{\alpha+l-\mu}}-\sum_{k=2}^{n}\mathrm{Li}_{\beta}\!\left(\frac{1}{k}\right)k^{\mu-\alpha}\right\},$
which serves as a natural extension of the Euler–Mascheroni constant. Here $\alpha,\beta\in\mathbb{C}$ with $\Re(\alpha)>0$ and $\mu\in\mathbb{N}_{0}$, and $\mathrm{Li}_{s}(z)=\sum_{l=1}^{\infty}z^{l}/l^{s}$ denotes the polylogarithm. When $\mu=0$, the corresponding sum is omitted. For $\mu\geqslant 1$,
(1.3) $D(\alpha,\beta;\mu)=\sum_{l=1}^{\mu}\frac{1}{l^{\beta}}-\gamma(\alpha,\beta;\mu).$
This paper establishes fundamental properties of the function $D(s_{1},s_{2};\lambda)$, including integral representations (Theorem 2.1 and Lemma 4.1), and provides closed-form expressions for its special values (Theorems 2.2–2.4 and 2.6) in terms of values of the Riemann zeta function and certain classical arithmetic constants. These results place $D(s_{1},s_{2};\lambda)$ within a unified framework that both incorporates earlier studies and reveals structural features that have not been explicitly formulated before.
Historically, Dirichlet series of the form $(1.1)$ originate in classical results of Goldbach, communicated in a letter to D. Bernoulli. In terms of the Riemann zeta function, Goldbach’s theorem may be written as
(1.4) $\sum_{k=2}^{\infty}\bigl\{\zeta(k)-1\bigr\}=1,$
which corresponds to the case $D(1,0;1)=1$. For $s_{1}=1$, $s_{2}=1$, and $\lambda=1$, identity $(1.3)$ reduces to Euler’s classical formula
@@
Following these pioneering works, numerous related formulas were obtained (see, for example, [Gl, Gl2, Raj, Ras, Ap]). A comprehensive historical account of closed-form evaluations of series involving the Riemann zeta function can be found in [Sr] and [SC, Chapter 3]. Motivated by these developments, it is natural to investigate explicit formulas for the special values of $D(s_{1},s_{2};\lambda)$. In the present paper, our approach relies on formulas due to Srivastava [Sr, (5.3)], Adamchik–Srivastava [AS, Proposition 4], and Choi–Srivastava [CS, (4.40)], which will be recalled in Section 2.
@@ 2. Main Results
Throughout this paper, $\Gamma(s)$ and $\zeta(s,\alpha)$ denote the gamma function and the Hurwitz zeta function, respectively. The symbol $\int_{+\infty}^{(0+)}$ denotes integration along a Hankel contour, which starts at positive infinity on the real axis, encircles the origin counterclockwise with a small radius, and then returns to the starting point. The generalized harmonic numbers are defined by $H_{m}^{(n)}:=\sum_{k=1}^{m}k^{-n}$, with $H_{m}:=H_{m}^{(1)}$ denoting the ordinary harmonic numbers, and we set $H_{0}^{(n)}:=0$. We write $B_{j}$ for the $j$th Bernoulli number. The binomial coefficient $\binom{x}{k}$ is defined by $x(x-1)\cdots(x-k+1)/k!$ for $k\in\mathbb{N}_{0}$.
@@
Let $\lambda\in\mathbb{C}$ with $\lambda\notin\mathbb{Z}_{<0}$. The Dirichlet series $D(s_{1},s_{2};\lambda)$ defined in (1.1) converges absolutely for all $s_{1},s_{2}\in\mathbb{C}$ except when $s_{1}\in\{0,-1,-2,\dots\}$. Moreover, it admits the following integral representation:
(2.1) $D(s_{1},s_{2};\lambda)=\frac{e^{-\pi is_{2}}\Gamma(1-s_{2})}{2\pi i\bigl(e^{2\pi is_{1}}-1\bigr)}\int\limits_{+\infty}^{(0+)}\int\limits_{+\infty}^{(0+)}\frac{v^{s_{1}-1}w^{s_{2}-1}}{e^{v+\lambda w}(e^{v}-1)}\sum_{k=1}^{\infty}\frac{(v/e^{w})^{k}}{\Gamma(s_{1}+k)}\,dv\,dw.$
This representation also defines a holomorphic function on the $(s_{1},s_{2})$-plane, except for $s_{1}\in\{0,-1,-2,\dots\}$.
@@
Let $\gamma(\alpha,\beta;\mu)$ be the generalized Euler constant defined in (1.2). For $n\in\mathbb{N}$ and $m,\mu\in\mathbb{N}_{0}$, one has
@@
Since classical times, it has been recognized that the arithmetic nature of Euler’s constant remains unresolved: its transcendence and the existence of any closed-form representation are unknown, and even its irrationality has not been established (for a comprehensive survey, see [Rag]). Likewise, a closed-form expression for $\gamma(n,m;\mu)$ $(n,m\in\mathbb{N},\ \mu\in\mathbb{N}_{0})$ also appears to be intractable. As shown in (1.5), the classical case $D(1,1;1)$ was already evaluated by Euler [Eu]. Furthermore, for $\mu=2,3$, the corresponding values $D(1,1;\mu)$ were obtained by Srivastava [Sr, (5.3)] and Choi–Srivastava [CS, (4.40)], in terms of the Stirling and Glaisher–Kinkelin constants, respectively:
$\begin{array}[]{l}D(1,1;2)={3}/{2}-{\gamma}/{2}-\log A_{0},\\[4.0pt] D(1,1;3)={11}/{6}-{\gamma}/{3}-\log A_{0}-2\log A_{1}.\end{array}$
Here $A_{0}=\sqrt{2\pi}$ is the Stirling constant, and $A_{1}$ is the Glaisher–Kinkelin constant, defined as follows:
$\log A_{0}$ $=\log\sqrt{2\pi}=\lim_{n\to\infty}\left\{\sum_{k=1}^{n}\log k-\left(n+\frac{1}{2}\right)\log n+n\right\},$
$\log A_{1}$ $=\lim_{n\to\infty}\left\{\sum_{k=1}^{n}k\log k-\left(\frac{n^{2}}{2}+\frac{n}{2}+\frac{1}{12}\right)\log n+\frac{n^{2}}{4}\right\}.$
The Glaisher–Kinkelin constant and its generalizations originally arose in the study of the hyperfactorial function and its extensions. In general, the constants $A_{m}$ ($m\in\mathbb{N}_{\geqslant 2}$) are referred to as the Bendersky constants (cf. [Be]; see also [Ad, Section 5], [Co], and [PG]). They are defined by the limit
$\log A_{m}:=\lim_{n\to\infty}\left\{\sum_{k=1}^{n}k^{m}\log k-p(n,m)\right\},$
@@
(2.3) $p(n,m)={}$ $\left(\frac{n^{m+1}}{m+1}+\frac{n^{m}}{2}\right)\log n-\frac{n^{m+1}}{(m+1)^{2}}$
$\quad+m!\sum_{r=1}^{m}\frac{B_{r+1}}{(r+1)!}\frac{n^{m-r}}{(m-r)!}\left(\log n+(1-\delta_{m,r})\sum_{i=1}^{r}\frac{1}{m-i+1}\right),$
with $\delta_{m,r}$ denoting the Kronecker delta. When $m=0$, the summation over $r$ is understood to be omitted. With these preliminaries in place, we are now ready to state one of our main theorems.
Theorem 2.3 (A relation between generalized Euler constants and the Bendersky constants).
@@
(2.4) $\gamma(1,1;\mu)=\frac{1}{\mu}\gamma+\sum_{j=0}^{\mu-2}\binom{\mu-1}{j}\log A_{j}.$
Consequently,
$D(1,1;\mu)=H_{\mu}-\frac{1}{\mu}\gamma-\sum_{j=0}^{\mu-2}\binom{\mu-1}{j}\log A_{j}.$
Here the summation over $j$ is understood to be empty when $\mu=1$.
@@
$\log A_{-1}(a):=\lim_{n\to\infty}\left\{\sum_{k=1}^{n}\frac{\log k}{k+a}-\frac{1}{2}\log^{2}n\right\}.$
Then
$D(1,1;0)=\sum_{l=1}^{\infty}(-1)^{l}\zeta^{\prime}(l+1)=\log A_{-1}(0)-\log A_{-1}(1).$
Equivalently,
$A_{-1}(a)=\lim_{n\to\infty}\prod_{k=1}^{n}\frac{k^{1/(k+a)}}{n^{\log\sqrt{n}}},\qquad e^{D(1,1;0)}=\prod_{n=1}^{\infty}n^{\frac{1}{n}-\frac{1}{n+1}}.$
In the above theorem, $\log A_{-1}(0)$ coincides with the first Stieltjes constant $\gamma_{1}$ (see, e.g., [Er, 1.12 (17)]). Theorems 2.3 and 2.4 show that the case $m=1$ is exceptional. In contrast, obtaining comparably simple closed-form expressions for $D(n,m;\mu)$ with $m\geqslant 2$ seems to require the introduction of a broader class of arithmetic constants. Nevertheless, one can derive the following identities.
@@
$\sum_{m=0}^{\infty}D(n,m;\mu)$ $=n+H_{\mu-1}-\gamma(n,1;\mu-1)-\sum_{k=2}^{n}\zeta(k),$
$\sum_{m=0}^{\infty}(-1)^{m}D(n,m;\mu)$ $=n-H_{\mu+1}+\gamma(n,1;\mu+1)-\sum_{k=2}^{n}\zeta(k),$
$\sum_{m=0}^{\infty}D(n,2m+1;\mu)$ $=H_{\mu}-\frac{1}{2}\left\{\gamma(n,1;\mu+1)+\gamma(n,1;\mu-1)+\frac{1}{\mu(\mu+1)}\right\}.$
@@
We now turn to the case $D(n,-m;\mu)$. In the fundamental case $n=\mu=1$, Adamchik and Srivastava obtained the following explicit formula [AS, Proposition 4] (see also [SC, Proposition 3.8]).
@@
Here $S(m,k)$ denotes the Stirling numbers of the second kind, defined by
@@
Building on this established result, we obtain the following theorem for $D(n,-m;\mu)$, which reduces to the above proposition in the case $n=\mu=1$.
@@
$D(n,-m;\mu)=\sum\limits_{l=0}^{m}\binom{m}{l}(\mu-n)^{m-l}D(1,-l;1)-\sum\limits_{k=2}^{n}\left\{\zeta(k)-1\right\}(k+\mu-n)^{m}.$
We also note that the following index-shifting relations for $D(n,\pm m;\mu)$ follow directly from Definition (1.1). Precisely, for $n,\mu\in\mathbb{N}$ and $m\in\mathbb{Z}\setminus\{0\}$:
$D(n,m;\mu)=\begin{cases}D(n-\mu+1,m;1)-\displaystyle\sum_{k=2}^{\mu}\frac{\zeta(k+n-\mu)-1}{k^{m}},&\text{if $n\geqslant\mu$,}\\[10.0pt] D(1,m;\mu-n+1)-\displaystyle\sum_{k=2}^{n}\frac{\zeta(k)-1}{(k+\mu-n)^{m}},&\text{if $n<\mu$.}\end{cases}$
By combining these index-shifting relations with Theorems 2.3 and 2.4, we obtain several structural relations among the values of $D(n,1;\mu)$.
As a concluding result of this section, we establish a functional relation for $D(s_{1},s_{2};\lambda)$ with respect to the variable $s_{2}$. In particular, for the values D(n,m;0)D(n,-m;0) with $n,m\in\mathbb{N}$ and $n>1$, this relation yields an explicit representation.
@@
Let $s_{1},s_{2}\in\mathbb{C}$ with $s_{1}\notin\mathbb{Z}_{\leqslant 0}$. Assume that $\Re(s_{1}-\lambda)>1$, $\Re(s_{2})<0$, and $1<\lambda\leqslant 0$. Then the following functional relation holds:
$D(s_{1},s_{2};\lambda)=\sum_{\begin{subarray}{c}k\geqslant 2,\,l\in\mathbb{Z}\end{subarray}}\frac{e^{2\pi il\lambda}\,\Gamma(1-s_{2})}{k^{\,s_{1}-\lambda}\,(2\pi il-\log k)^{\,1-s_{2}}}.$